Three defective electric toothbrushes were accidently shipped along with 17 non-defective.
a. what is the probability that the first two selected will be defective.
b. what is the probability the first two sold will not be defected

Answer by stanbon(75887)

Three defective electric toothbrushes were accidently shipped along with 17 non-defective.
a. what is the probability that the first two selected will be defective.
Ans: 3C2/17C2 = 3/136
------------------------------------------
b. what is the probability the first two sold will not be defected
Ans: 14C2/17C2 = 0.669...
Cheers,
Stan H.